A man is walking over a dome of 10 m in radius. how far can he descend from the dome without slipping? take the coefficient of friction between the surface of the dome and the shoe of the man 0.6.


1
Expert's answer
2021-03-10T17:16:56-0500

Explanations & Calculations


  • Refer to the sketch attached
  • Consider a situation when the person is at an arbitrary point on the dome as shown & write equations for his equilibrium.


  • The contact force on him at that point from the dome

"\qquad\qquad
\begin{aligned}
\small R&= \small mg\cos\theta
\end{aligned}"

  • The person can safely descend some distance down until the component of his weight becomes equal to the maximum static friction available at that point.


  • Then,

"\qquad\qquad
\begin{aligned}
\small f_{s.max}&\geq\small mg\sin\theta\\
\small \mu R&\geq \small mg\sin\theta\\
\small \mu\cdot mg\cos\theta&\geq\small mg\sin\theta \\
\small \tan\theta&\leq\small \mu\\
\theta&\leq\small \tan^{-1}(0.6)\\
&\leq\small 30.96^0
\end{aligned}"

  • As mentioned above he can safely come down up to a maximum angle of 30.96 degrees from the verticle
  • Then

"\qquad\qquad
\begin{aligned}
\small \theta _{max}&= \small 30.96^0=0.540rad
\end{aligned}"

  • The distance he can come down is (distance measured from the top of the dome)

"\qquad\qquad
\begin{aligned}
\small S&= \small r\theta\\
&= \small 10m\times 0.540\\
&= \small \bold{5.4\,m}
\end{aligned}"